# The Representation and Recall of Interwoven Structured Knowledge in LLMs: A Geometric and Layered Analysis

### v1

@@ Abstract
This study investigates how large language models (LLMs) represent and recall multi-associated attributes across transformer layers. We show that intermediate layers encode factual knowledge by superimposing related attributes in overlapping spaces, along with effective recall even when attributes are not explicitly prompted. In contrast, later layers refine linguistic patterns and progressively separate attribute representations, optimizing task-specific outputs while appropriately narrowing attribute recall. We identify diverse encoding patterns including, for the first time, the observation of 3D spiral structures when exploring information related to the periodic table of elements. Our findings reveal a dynamic transition in attribute representations across layers, contributing to mechanistic interpretability and providing insights for understanding how LLMs handle complex, interrelated knowledge.
@@ 1 Introduction
Large language models (LLMs) excel in tasks like translation, text generation, and comprehension 27, 5, 18, yet their mechanisms for storing and processing information remain poorly understood. In human cognition, complex networks of interrelated knowledge enable efficient generalization, inference, and creativity by integrating new information into existing frameworks 21, 3. But how do LLMs represent multi-related attributes and interconnected knowledge? For example, when knowing a chemical element like hydrogen, are attributes such as ‘the lightest element’ and ‘atomic number 1’ stored independently, or do they influence each other? When one attribute is activated, does it trigger related concepts through a network of connections? If these attributes interact, how are they organized—linearly, cyclically, or in more complex structures? Clarifying these mechanisms is vital for aligning LLMs with human values 16, enhancing their design, and broadening their applications.
Mechanistic interpretability offers a pathway to answer these questions, aiming to reverse-engineer neural networks into human-understandable algorithms 4, 30, 6. A proposed research direction is the linear representation theory, suggesting that LLMs encode knowledge as one-dimensional lines, with model states formed by sparse combinations of these representations. Evidence supports this for spatial, temporal, and sentiment knowledge 12, 33. However, emerging evidence points to more intricate structures, such as circular encodings for periodic concepts like days or months 10, challenging the simplicity of the linear model.
Building on prior work that primarily examined single-attribute representations, we delve deeper into how LLMs encode and recall complex, interwoven knowledge. Our study explores the interaction and independence of linguistic and factual representations, as well as their underlying non-linear geometric relationships. The key findings of our study are:
1. In LLMs, intermediate layers tend to focus on factual knowledge, while later layers shift toward linguistic patterns and task-specific outputs (Sec.3).
2. LLMs can recall associative knowledge about related attributes, even when not explicitly mentioned in the prompt, with the strongest recall observed in intermediate layers, diminishing in deeper layers (Sec.4.1).
3. Recall mechanisms exhibit diverse geometric relationships. Notably, we identify 3D spiral structures encoding interrelated attributes for the first time (Sec.4.2).
4. Attribute representations are superimposed in intermediate layers, exhibiting linear relationships. Later layers focus on refining these representations, leading to clearer separation (Sec.5).
@@ 2 Preliminaries
Our study only focuses on how reliably acquired knowledge (i.e. things we’re confident the model knows) is represented within LLMs, and excludes hallucinations or information not in the training set. We use the properties of chemical elements in the periodic table as a case study due to their frequent occurrence in training data, well-defined attributes, and quantifiable properties, making them an ideal subject for this investigation. We adopt Llama series models in this study.
2.1 Activation collection
Generating prompts
@@ 2.1 Activation collection
Each prompt $p_{i,j,k}$ will then be fed into LLMs to study the corresponding activations at different layers.
Collecting last-token activations
Last-token activations capture the full prompt context in decoder-only models with masked attention, as they integrate information from all preceding tokens. For each layer $l$, we collect last-token activations $\mathbf{h}_{i,j,k}^{(l)}$ from prompts $p_{i,j,k}$ across all elements and templates:
where $f^{(l)}$ denotes the layer-$l$ transformation, $T$ is the token length, and $d$ is the hidden dimension. The initial activation $\mathbf{h}_{i,j,k}^{(0)}$ is obtained by embedding the prompt through an embedding layer $E_{0}$, followed by processing through $L$ Transformer layers. Each layer applies multi-head attention and a feedforward network with residual connections and layer normalization:
$\mathbf{h}^{\prime(l)}_{i,j,k}=\text{LayerNorm}\left(\mathbf{h}^{(l-1)}_{i,j,k}+\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})\right)$
$\mathbf{h}^{(l)}_{i,j,k}=\text{LayerNorm}\left(\mathbf{h}^{\prime(l)}_{i,j,k}+\text{FFN}(\mathbf{h}^{\prime(l)}_{i,j,k})\right).$
$\text{logits}_{i,j,k}=\mathbf{h}_{i,j,k}^{(L)}\mathbf{W}_{\text{vocab}}$
By analyzing last-token activations $\mathbf{h}_{i,j,k}^{(l)}$ across layers, we investigate how attributes are represented in the model’s hidden states.
2.2 Activation distribution
We start with a preliminary visualization of the distribution of last-token activations for the ‘atomic number’ attribute. Activations from each transformer layer $l$ were collected for the atomic number attribute across the first 50 elements using 11 continuation-style templates, forming the set $H_{\text{atomic number}}^{(l)}$. To enable informative plots to be produced efficiently, PCA was applied to $H_{\text{atomic number}}^{(l)}$ and then t-SNE was use to project the first 50 principle components into 2D. Fig.1 shows the resulting distributions for Meta-Llama-3.1-70B, with points colored by atomic number and other attributes, revealing their associations to atomic number.
The first column of the figure colors activations by true atomic number values (explicitly mentioned in the prompt). In early layers, prompts with similar vocabulary cluster together irrespective of atomic number, reflecting token-level similarity. In the intermediate layer, the activations are hierarchically clustered. The small clusters each contain activations for the 11 prompts generated for each single element. These small clusters then exhibit clustering with respect to the other elements, showing that the model encoding attribute-specific knowledge. By the final layers, prompts with the same meaning remain clustered, but distinctions between them become clearer, reflecting refined language and factual understanding.
Figure 1: t-SNE visualization of Meta-Llama-3.1-70B last-token activations from the 1st, 50th, and 80th layers, using 11 continuation-style templates across 50 elements (550 points per plot). Each column shows one layer, while rows represent different colormaps highlighting attributes: ‘atomic number’, ‘group’, ‘period’, and ‘category’. In the top-left plot, circled clusters correspond to individual templates, each containing 50 points.
In the next three columns, the activations are colored by the true values of attributes unmentioned in the prompt: ‘group’, ‘period’, and ‘category’. Despite not being mentioned in the prompt, all attributes cluster in the intermediate layers. By the final layers, the clustering of some attributes, such as group and period, becomes less coherent, indicating a shift in representation.
@@ 2.2 Activation distribution
Based on the observed activation distributions, we hypothesize that intermediate layers in LLMs focus on factual knowledge, while later layers refine linguistic patterns and task-specific outputs, with attributes transitioning from being superimposed in intermediate layers to distinct in later layers. Additionally, attributes are encoded in geometric patterns reflecting their properties. These hypotheses are validated in the following sections.
3 Intermediate layers encode knowledge, later layers shape language
Figure 2: $R^{2}$ score trends across layers for different models and attributes. Linear probing with 5-fold cross-validation was performed on activations, and $R^{2}$ scores on the test set are shown for each attribute.
To understand how complex interrelated knowledge emerges, we first investigate how individual attributes are represented and evolve across layers. As a preliminary step, we analyzed attention maps, revealing that intermediate layers focus more on tokens with a significant impact on the output, while later layers distribute attention more evenly, suggesting a transition from capturing specific relationships to integrating broader context for cohesive outputs. Detailed findings are provided in the appendix (see Fig.A1). The following sections quantitatively examine the roles of intermediate and later layers in representing attributes within LLMs.
3.1 Intermediate layers know factual knowledge. Language? Not yet!
We used linear probing, with 5-fold cross-validation, to train linear Support Vector Regression (SVR) models for each layer, $l$, and attribute, $A_{j}$, based on the activation dataset $H_{j}(l)$. The ground truths correspond to attribute values $y_{i,j}$ for each element $X_{i}$. To focus on numerical factual knowledge while minimizing the influence of language patterns, we used continuation-style prompts. Specifically, SVR maps each activation vector $\mathbf{h}_{i,j}^{(l)}$ from layer $l$ to the target attribute $y_{i,j}$ using the following linear function:
$\hat{y}_{i,j}^{(l)}=\mathbf{w}_{j}^{(l)}\cdot\mathbf{h}_{i,j}^{(l)}+b_{j}^{(l)},$
where $\mathbf{w}_{j}^{(l)}\in\mathbb{R}^{d}$ is the weight and $b_{j}^{(l)}\in\mathbb{R}$ is the bias.
The $R^{2}$ score trends of the test set for each attribute across layers are shown in Fig.2 with the detailed results for the best layer provided in Appendix D.2. Both the intermediate and last layers exhibit high $R^{2}$ scores (note that is not 1; see Appendix D.1). The last layer’s performance is expected, as it generates the final outputs. The intermediate layers’ high $R^{2}$ scores indicate that factual and numerical knowledge is already effectively encoded at this stage, suggesting LLMs represent complex knowledge earlier than expected. These findings also explain the attention map results, where intermediate layers focus on key tokens because they have already encoded conceptual knowledge at this stage.
Intermediate layers can represent factual knowledge, but can they also express this knowledge in human language? We input the prompt ‘The atomic number of Mg is ’ and analyzed the token probabilities at each layer. By normalizing the final token’s hidden state with LayerNorm and applying the vocabulary head followed by softmax, we obtained the top-ranked tokens directly output by each layer ***Applying language heads to the hidden states of intermediate layers, known as early exit  32, 3, 9, has proven effective even without extra training 19.. In each layer, we extracted the probability of the target token, $t_{\text{target}}$ — the output token from the last layer, and checked if it ranked within the top 50 most probable tokens for that layer. The results are shown in Fig.3.
Figure 3: Probability of the target token $t_{\text{target}}$ across layers in Meta-Llama-3-8B for the prompt ‘The atomic number of Mg is ’ Each line shows $t_{\text{target}}$’s probability at each layer. The probabilities were calculated by iteratively re-running the model with the next token added to the prompt. Markers indicate the layers where it ranks in the top 50 most probable tokens.
3.2 Later layers shift focus from factual knowledge to language patterns
The previous experiment shows that later layers begin to transform factual knowledge into language patterns. To further validate this, we compared question-style and continuation-style prompts using linear probing, following the procedure in Sec.3.1. The attribute ‘group’ was used as an example, with results shown in Fig.4.
Figure 4: $R^{2}$ Scores for ‘Group’ Attribute: Linear probing comparison of continuation and question prompts across layers.
The results show a significant drop in $R^{2}$ scores for question prompts compared to continuation prompts in the later layers, with the gap widening in deeper layers. While both prompt types reflect a shift toward language structure in the later layers, continuation prompts are less affected because their output remains tied to factual tokens. In contrast, question prompts emphasize sentence structure and may generate filler words like ‘The,’ leading to the intentional discarding of factual representation. This effect is more pronounced in smaller models, while larger models retain more information about factual knowledge while processing language.
4 Recall peaks at intermediate layers
The previous section focused on the behavior of single attributes across layers. Here, we investigate whether related attributes are interconnected by examining the recall ability of LLMs—their capacity to retrieve attributes related to, but not explicitly mentioned in the prompt. Additionally, we analyze the geometric mechanisms underlying this recall process.
4.1 Recalling knowledge ability peaks at intermediate layers and declines afterward
We conducted an experiment to explore the relationship between different attributes using misaligned linear probing. Specifically, we generated activation datasets from prompts about $H_{\text{atomic number}}^{(l)}$ and $H_{\text{group}}^{(l)}$. Separate probes were trained on each activation dataset, but in both cases, only $y_{\text{group}}$ (group labels) were used as the target during training. This approach examines whether activations related to one attribute (e.g., atomic number) encode information about another attribute (e.g., group). Continuation prompts were used, and the probing process followed the method described in Sec.3.1. The results are shown in Fig.5.
The results showed that in the early and intermediate layers, $R^{2}$ scores for recalling group information were consistent, regardless of whether activations came from prompts about atomic number or group. However, in the later layers, activations from atomic number prompts showed a significant drop in $R^{2}$ scores for group recall compared to those from group prompts. This suggests that intermediate layers enable LLMs to recall related attribute knowledge, indicating a capacity for broader knowledge representation. The significant drop in the later layers suggests that knowledge not immediately necessary for generating the next token is not activated. The relationships between different attribute representations are further explored in Sec.5.
Figure 5: $R^{2}$ Score Trends for Misaligned Linear Probing. The probe was trained on ‘group’ values using activations from prompts about ‘group’ (matching) and ‘atomic number’ (non-matching).
4.2 Attribute geometric interrelationship
Figure 6: Activation Patching Results for Layer 20 in Meta-Llama-3.1-70B. The model’s predictions are evaluated after replacing the activation of the ‘element’ token at the last token position with the predicted activation $\hat{\mathbf{h}}^{\text{pred},(20)}_{0}$.
@@ 4.2 Attribute geometric interrelationship
LLMs can recall knowledge in the early to intermediate layers, but how do these attributes interact? We hypothesize that attributes in LLMs exist in a high-dimensional space, manifesting as linear, circular, or spiral patterns based on their structure, and then proceed to validate these geometries.
@@
Inspired by 10, we map the last-token activations $\mathbf{h}^{(l)}\in\mathbb{R}^{k}$ at layer $l$ to a geometric space $f(r,g,p)$, which represents a geometric structure defined by atomic number $r$, group $g$, and period $p$. To achieve this, we first reduce the dimensionality of the activations using PCA, denoted as $\mathbf{P}(\mathbf{h}^{(l)})$, followed by layer-wise linear regression:
$\mathbf{W}^{(l)},\mathbf{b}^{(l)}=\arg\min_{\mathbf{W}^{\prime},\mathbf{b}^{\prime}}\sum_{i}\left\|\mathbf{W}^{\prime}\mathbf{P}(\mathbf{h}^{(l)}_{i})+\mathbf{b}^{\prime}-f_{i}\right\|_{2}^{2}$
@@
Here, $\mathbf{W}^{(l)}\in\mathbb{R}^{d^{\prime}\times k}$ and $\mathbf{b}^{(l)}\in\mathbb{R}^{d^{\prime}}$ are the learned weight and bias for layer $l$. $f_{i}=f(r_{i},g_{i},p_{i})$ denotes the mapping of the $i$-th element in the geometric space.
To perform an intervention, we calculate the centroid of the PCA-reduced activations for layer $l$, denoted as $\mathbf{\bar{h}}^{(l)}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{P}(\mathbf{h}^{(l)}_{i})$, and map it to the geometric space using $\mathbf{W}^{(l)}\mathbf{\bar{h}}^{(l)}+\mathbf{b}^{(l)}$. The mapping of the target element is $f_{0}=f(r_{0},g_{0},p_{0})$. The deviation from this mapping is then mapped back to the activation space using the pseudo-inverse of the learned weight, $(\mathbf{W}^{(l)})^{+}$, yielding the predicted activation of the target element:
$\hat{\mathbf{h}}^{\text{pred},(l)}_{0}=\mathbf{P}^{-1}\left(\mathbf{\bar{h}}^{(l)}+(\mathbf{W}^{(l)})^{+}\left(f_{0}-(\mathbf{W}^{(l)}\mathbf{\bar{h}}^{(l)}+\mathbf{b}^{(l)})\right)\right)$
Importantly, the model never accesses the original activation of the target element; the predicted activation is solely based on the mapping of target element in the geometric space and the activations of other elements. During inference, we replace the activation of the ‘element’ token (last token position) in the 20th layerDetails of intervention performance are provided in Appendix B.1. From layer 20, the interventions show effectiveness. with $\hat{\mathbf{h}}^{\text{pred},(20)}_{0}$, using the prompt ‘In the periodic table, the atomic number of element’ The results are evaluated to verify if the model predicts the target properties without using the original activation.
@@
Figure 7: Predicted Atomic Numbers After Intervention on Difference Geometric Space. Left: $(\cos\theta,\sin\theta,r)$. Right: $(r\cos\theta,r\sin\theta,r)$. The colored points represent the tokens with the highest logits output by the model after intervention.
@@
We evaluate the effectiveness of different geometric spaces for interventions, including linear, 2D spiral, and 3D spiral geometries. Angular variables $\theta=\frac{2\pi g}{18}$ are used to capture periodic relationships. To test the impact of disrupted geometry, two random spaces are introduced: in Space 8, atomic numbers $r$ are shuffled; in Space 9, $\theta$ is randomly permuted. Additionally, in Space 10, the prompt ‘In numbers, the Arabic numeral for number’ generates numbers 1–50, testing whether periodic patterns emerge without explicit element references. Results are shown in Fig.6, with detailed values in Table B1 in appendix.
@@
Effective activation patching suggests that the target space $f(r,g,p)$: 1) retains sufficient information for accurate reconstruction during transformations with the activation space, and 2) preserves geometric structures similar to those in the activation space to ensure valid adjustments in the high-dimensional space.
Results show that intervention can be applied in various geometric spaces, with some performing significantly better. Spaces such as $(\cos\theta,\sin\theta,r)$ and $(r\cos\theta,r\sin\theta,r)$ over $70\%$ predictions of the atomic number have an absolute error within 2, suggesting the potential existence of latent 3D structures in LLMs resembling spirals or radial spirals. Fig.7 shows the output of the LLM after intervention on the two spiral geometric space, highlighting the periodic table’s cyclic properties. Additional geometric shapes are analyzed in Appendix B.2. In contrast, random spaces and unrelated prompts exhibit poor performance, increasing our confidence that these results suggest the geometry of embedding spaces can relate to the real geometry of knowledge.
@@
Figure 8: Euclidean distance heatmap of approximated vector representations for numeric tokens (1-50) in the hidden space of the last layer.

### v2

# Layerwise Recall and the Geometry of Interwoven Knowledge in LLMs

@@ Abstract
This study explores how large language models (LLMs) encode interwoven scientific knowledge, using chemical elements and LLaMA-series models as a case study. We identify a 3D spiral structure in the hidden states that aligns with the conceptual structure of the periodic table, suggesting that LLMs can reflect the geometric organization of scientific concepts learned from text. Linear probing reveals that middle layers encode continuous, overlapping attributes that enable indirect recall, while deeper layers sharpen categorical distinctions and incorporate linguistic context. These findings suggest that LLMs represent symbolic knowledge not as isolated facts, but as structured geometric manifolds that intertwine semantic information across layers. We hope this work inspires further exploration of how LLMs represent and reason about scientific knowledge, particularly in domains such as materials science.
@@ 1 Introduction
Large language models (LLMs) have demonstrated strong factual recall across a wide range of domains, from history and geography to science and mathematics (Naveed et al., 2023; Chang et al., 2024; Kaddour et al., 2023). However, much remains unknown about how interconnected knowledge is internally organized. Clarifying these mechanisms is vital for aligning LLMs with human values (Ji et al., 2023), enhancing their design, and broadening their applications.
Mechanistic interpretability offers a pathway to answer these questions (Bereska and Gavves, 2024; Singh et al., 2024; Dar et al., 2022). Prior work suggests that models often represent features as linear directions in activation space (Gurnee and Tegmark, 2023; Tigges et al., 2023). However, growing evidence reveals non-linear and geometrically structured patterns—such as cycles and modular circuits in tasks involving periodic concepts like days and months(Engels et al., 2024). These observations raise a deeper question: can LLMs represent interwoven knowledge in forms that mirror more complex conceptual geometry in the real world?
Prior studies on factual recall have shown that core facts are enriched in early and middle multilayer perceptron (MLP) layers, propagated by attention, and linearly decodable via circuit-level analysis(Meng et al., 2022; Geva et al., 2023; Nanda et al., 2023b). They largely focus on isolated attributes, leaving open how multiple, interrelated facts are jointly encoded and retrieved.
Building on this, we delve into how LLMs encode and recall complex, interwoven knowledge through global, structured, and geometry-aware representations that evolve across the model’s depth. The contribution of our study are:
1. We report the first observation of a 3D spiral structure in LLM hidden states that organizes chemical elements in alignment with the structure of periodic table (Sec 3).
2. We compare regression and classification probing, showing that middle layers encode continuous attribute structure, while later layers sharpen boundaries for fine-grained decisions (Sec 4.1).
3. We show that linguistic structure increasingly shapes knowledge representations in later layers (Sec 4.2).
4. We find that LLMs recall related attributes through strong linear associations in middle layers, which weaken in deeper layers (Sec 5).
Figure 1: t-SNE visualization of Meta-Llama-3.1-70B last-token residual streams from the 1st, 50th, and 80th layers, using 11 continuation-style templates across 50 elements (550 points per plot). Each column shows one layer, while rows represent different colormaps highlighting attributes: ‘atomic number’, ‘group’, ‘period’, and ‘category’. In the top-left plot, circled clusters correspond to individual templates, each containing 50 points.
@@ 2 Preliminaries
Our study only focuses on how reliably acquired knowledge (i.e. things we’re confident the model knows) is represented within LLMs, and excludes hallucinations or information not in the training set. We use the properties of chemical elements in the periodic table as a case study due to their frequent occurrence in training data, well-defined attributes, and quantifiable properties, making them an ideal subject for this investigation. We adopt Llama series models (Touvron et al., 2023; Grattafiori et al., 2024) in this study.
2.1 Residual stream collection
@@
Each prompt $p_{i,j,k}$ can then be fed into LLMs to study the corresponding residual streams at different layers.
Last-token residual streams capture the full prompt context in decoder-only models with masked attention, as they integrate information from all preceding tokens. For each layer $l$, we collect last-token residual streams $\mathbf{h}_{i,j,k}^{(l)}$ from prompts $p_{i,j,k}$ across all elements and templates (see Appendix A for details).
2.2 Residual stream distribution
We start with a preliminary visualization of the distribution of last-token residual streams for the ‘atomic number’ attribute. Residual streams from each transformer layer $l$ were collected for the atomic number attribute across the first 50 elements using 11 continuation-style templates, forming the set $H_{\mathrm{AtomicNumber}}^{(l)}$. To enable informative plots to be produced efficiently, PCA was applied to them and then t-SNE was use to project the first 50 principle components into 2D. Fig.1 shows the resulting distributions for Meta-Llama-3.1-70B, with points colored by atomic number as well as the other attributes (to visualize their association to atomic number).
The first column of the figure colors residual streams by true atomic number values (explicitly mentioned in the prompt). In early layers, prompts with similar vocabulary cluster together irrespective of atomic number, reflecting token-level similarity. In the middle layers, residual streams for the same element form small, heavily overlapping clusters, each containing representations from the 11 prompts generated for that element. By the final layers, although these points (same colors) still cluster together, individual points become distinguishable due to increased spread within the clusters.
In the next three columns, the residual streams are colored by the true values of attributes unmentioned in the prompt: ‘group’, ‘period’, and ‘category’. Despite not being mentioned in the prompt, in middle layers, chemically similar elements (small clusters with similar colors) cluster closely together. By the final layers, the clustering of some attributes, such as group and category, becomes less coherent, indicating a shift in representation.
@@ 2.2 Residual stream distribution
These visualizations suggest that LLM residual streams may encode attribute relationships in a structured and potentially geometric manner that reflect properties of the physical world. In particular, intermediate layers appear to capture implicit similarities even for unmentioned attributes, while later layers are tuned for task-specific outputs. These hypotheses are examined in the following sections.
3 Geometric relationships among attributes
In materials science, a spiral trend emerges from the periodic variation in valence electron configurations as atomic number increases. By arranging elements sequentially and mapping their properties in a polar coordinate system, this periodicity becomes visually apparent as a spiral. We investigate whether LLMs, trained on extensive data about elemental properties, inherently capture these physical periodicities and reflect similar spiral structures in their learned embeddings.
@@ 3 Geometric relationships among attributes
We hypothesize that attributes in LLMs exist in a high-dimensional space, manifesting as linear, circular, or spiral patterns based on their structure, and then proceed to validate these geometries.
Figure 2: Residual Stream Patching Results for Layer 20 in Meta-Llama-3.1-70B. The model’s predictions are evaluated after replacing the residual stream of the ‘element’ token at the last token position with the predicted residual stream $\hat{\mathbf{h}}^{\text{pred},(20)}_{0}$.
@@
Inspired by Engels et al. (2024), we map the last-token residual streams $\mathbf{h}^{(l)}\in\mathbb{R}^{k}$ at layer $l$ to a geometric space $f(r,g,p)$, which encodes atomic number $r$, group $g$, and period $p$. To learn this mapping, we first reduce the dimensionality of the residual streams to 30 using PCA, denoted as $\mathbf{P}(\mathbf{h}^{(l)})$, and fit a linear projection using all elements except one held-out target:
$\mathbf{W}^{(l)},\mathbf{b}^{(l)}=\arg\min_{\mathbf{W}^{\prime},\mathbf{b}^{\prime}}\sum_{i\neq 0}\left\|\mathbf{W}^{\prime}\mathbf{P}(\mathbf{h}^{(l)}_{i})+\mathbf{b}^{\prime}-f_{i}\right\|_{2}^{2}$
@@
where $\mathbf{W}^{(l)}\in\mathbb{R}^{d^{\prime}\times 30}$, $\mathbf{b}^{(l)}\in\mathbb{R}^{d^{\prime}}$, and $f_{i}=f(r_{i},g_{i},p_{i})$ denotes the mapping of the $i$-th element in the geometric space.
To perform the intervention, we compute the centroid of the PCA-reduced residual streams for the remaining $N=K-1$ elements:
$\bar{\mathbf{h}}^{(l)}=\frac{1}{N}\sum_{i\neq 0}\mathbf{P}(\mathbf{h}^{(l)}_{i})$
then map it to the geometric space: $\mathbf{z}=\mathbf{W}^{(l)}\bar{\mathbf{h}}^{(l)}+\mathbf{b}^{(l)}$. Let $f_{0}=f(r_{0},g_{0},p_{0})$ denote the target element’s embedding in the geometric space. The deviation $f_{0}-\mathbf{z}$ is projected back to the residual stream space using the pseudo-inverse of $\mathbf{W}^{(l)}$, giving the predicted (intervened) residual stream:
$\hat{\mathbf{h}}^{\text{pred},(l)}_{0}=\mathbf{P}^{-1}\left(\bar{\mathbf{h}}^{(l)}+\left(\mathbf{W}^{(l)}\right)^{+}(f_{0}-\mathbf{z})\right)$
@@
Figure 3: Predicted atomic numbers after intervention in 3D spiral space $(r\cos\theta,r\sin\theta,r)$. Colored points indicate the tokens with highest logits.
Importantly, the model never accesses the original residual stream of the target element; the predicted residual stream is computed solely from its geometric representation and the residual streams of other elements. During inference, we replace the residual stream of ‘element’ (last token position) in the 20th layer [^1] See Appendix B.1 for details on intervention performance. Interventions become effective from layer 20 onward. with $\hat{\mathbf{h}}^{\text{pred},(20)}_{0}$, using the prompt ‘In the periodic table, the atomic number of element’. We then evaluate whether the model can correctly output the target token without ever seeing its original residual stream.
@@
We evaluate the effectiveness of different geometric spaces for interventions, including linear, 2D spiral, and 3D spiral geometries. Angular variables $\theta=\frac{2\pi g}{18}$ are used to capture periodic relationships. To test the impact of disrupted geometry, two random spaces are introduced: in Space 8, atomic numbers $r$ are shuffled; in Space 9, $\theta$ is randomly permuted. Additionally, in Space 10, the prompt ‘In numbers, the Arabic numeral for number’ generates numbers 1–50, testing whether periodic patterns emerge without explicit element references.
@@
Effective residual stream patching suggests that the target space $f(r,g,p)$: 1) retains sufficient information for accurate reconstruction during transformations with the residual stream space, and 2) preserves geometric structures similar to those in the residual stream space to ensure valid adjustments in the high-dimensional space.
Patching results for Meta-Llama-3.1-70B are shown in Fig. 2, with detailed values in Table B.1 (Appendix). Results show that intervention can be applied in various geometric spaces, with some performing significantly better. Spaces such as $(\cos\theta,\sin\theta,r)$ and $(r\cos\theta,r\sin\theta,r)$ over $70\%$ predictions of the atomic number have an absolute error within 2, suggesting the potential existence of latent 3D structures in LLMs resembling spirals or radial spirals. Fig. 3 illustrates the LLM’s output post-intervention in 3D spiral geometry. Additional geometric analyses are in Appendix B.2. Random spaces and unrelated prompts perform poorly, reinforcing the correlation between embedding geometry and real-world knowledge structures. In a concurrent study, Kantamneni and Tegmark (2025) observed spiral-like structures in number space with periods of 2, 5, 10, and 100, likely reflecting common human conventions in numerical representation. In contrast, our model exhibits a distinct 18-period spiral aligned with the periodic structure of chemical elements. This representation performs notably worse for ordinary numbers without elemental context (which aligns with their observation that the 18-period does not prominently emerge), indicating that such geometric patterns emerge from underlying physical or semantic regularities rather than arbitrary structures.
@@
Figure 4: Euclidean distance heatmap of approximated vector representations for numeric tokens (1-50) in the hidden space of the last layer in Meta-Llama-3.1-70B.